OpenStudy (anonymous):

Line Q is represented by the following equation: 2x + y = 11 Which equation completes the system that is satisfied by the solution (3, 5)? A. x + 2y = 15 B. x + y = 15 C. x + y = 8 D. x − y = 2

OpenStudy (anonymous):

@ayeshaafzal221

OpenStudy (anonymous):

would it be C?

OpenStudy (anonymous):

hw did u come up with this answer?

OpenStudy (anonymous):

did u solve each eq simultaneously, because when u do when u get x or y value same as (3,5) that equation will be ur answer

OpenStudy (anonymous):

btw ur answer is right

OpenStudy (anonymous):

I get how to substitute the numbers but i dont fully understand how to answer the question. Am i simplifying?

OpenStudy (anonymous):

no its a very long question basically u solve each eq simultaneously with original eq "2x+y=11"

OpenStudy (anonymous):

it would end up being 2x3 + 5 = 11 so how to the answer choices come up?

OpenStudy (anonymous):

2x+y=11 x+y=8 2x+y=11 -2(x+y=8) 2x+y=11 -2x-2y=-16 -y=-5 y=5

OpenStudy (anonymous):

like this ^

OpenStudy (anonymous):

but if u want shorter way u can sub ur value ur right

OpenStudy (anonymous):

ah, well thanks.